# Image Demoireing in RAW and sRGB Domains

### 14 Dec 2023 v1

@@ Abstract
Moiré patterns frequently appear when capturing screens with smartphones or cameras, potentially compromising image quality. Previous studies suggest that moiré pattern elimination in the RAW domain offers greater efficiency compared to demoiréing in the sRGB domain. Nevertheless, relying solely on raw data for image demoiréing is insufficient in mitigating color cast due to the absence of essential information required for color correction by the Image Signal Processor (ISP). In this paper, we propose perform Image Demoiréing concurrently utilizing both RAW and sRGB data (RRID), which is readily accessible in both smartphones and digital cameras. We develop Skip-Connection-based Demoiréing Module (SCDM) with specific modules embeded in skip-connections for the efficient and effective demoiréing of RAW and sRGB features, respectively. Subsequently, we propose RGB Guided Image Signal Processor (RGISP) to incorporate color information from coarsely demoiréd sRGB features during the ISP stage, assisting the process of color recovery. Extensive experiments demonstrate that our RRID outperforms state-of-the-art approaches by 0.62dB in PSNR and 0.003 in SSIM, exhibiting superior performance both in moiré pattern removal and color cast correction.
@@ 1 Introduction
With the increasing prevalence of smartphone cameras, it is common to record on-screen content using smartphone cameras for convenience. However, this type of photography often results in the presence of moiré patterns in the captured images. The appearance of moiré patterns is caused by spatial frequency aliasing between the camera’s color filter array (CFA) and the screen’s LCD subpixel, resulting in an unsatisfactory visual experience.
@@
Moiré patterns present challenges to be eliminated due to their different scales, indistinct shapes, diverse colors, and varying frequencies. Nowadays, numerous learning-based image demoiréing approaches [20, 2, 6, 7, 15, 32, 16, 22, 26, 17, 21] have been proposed. Among the various methods employed, a commonly used approach is utilizing a multi-scale architecture to remove various sizes of moiré patterns [20, 2, 6, 26, 17]. Besides, some works advocate the leverage of frequency components and delicately design frequency-based demoiréing networks [7, 15, 32, 21].
@@
The aforementioned methods are developed to remove moiré patterns in the standard RGB (sRGB) domain, as depicted in Fig. 1(a). While this scheme seems to be the most direct approach, its effectiveness is limited in removing moiré patterns from complex scenes. The reason for this limitation probably lies in the deterioration of moiré patterns in the RAW domain, caused by the nonlinear operation during the Image Signal Processor (ISP) process, such as demosaicing. Prior researches consequently suggest that eliminating moiré patterns in the RAW domain is more efficient than performing demosaicing in the sRGB domain [27, 29]. Due to the easy accessibility of RAW data in both smartphones and digital cameras, utilizing it for image demoiréing can be both feasible and advantageous. RDNet [27] is the pioneering study that investigates image demoiréing in the RAW domain. The researchers explore the demoiréing of RAW images using an encoder-decoder and class-specific learning approach, followed by the utilization of a pre-trained ISP for the RAW to sRGB domain conversion process, as depicted in Fig. 1(b). Nevertheless, relying exclusively on RAW data can lead to color cast in the restored images due to the absence of essential information in the ISP.
@@
In order to correct the color cast and make the best use of RAW and sRGB information for moiré pattern removal, we propose image demoiréing in both the RAW and sRGB domains (RRID). Firstly, we advocate that introducing RAW data is advantageous moiré removal due to the following reasons: (1) RAW pixels provide more information compared to sRGB pixels since they are typically 12 or 14 bits. (2) Moiré patterns in the RAW domain are less apparent because they are not further affected by nonlinear operations in the ISP. In addition, we further incorporate sRGB data as the color transformation guidance in the ISP, as demonstrated in Fig. 1(c), facilitating color deviations correction. In RRID, we adopt a multi-scale architecture in the Skip-Connection-based Demoiréing Module (SCDM), to effectively eliminate moiré patterns of different sizes. In order to efficiently perform demoiréing on both RAW data and sRGB features, distinct modules are integrated into skip-connections within SCDM, as opposed to brutally injecting them into the multi-scale layers. Specifically, for RAW data demoiréing, we introduce the Gated Feedback Module (GFM) into the skip connection, enabling adaptive differentiation between texture details and moiré patterns through feature gating. Moreover, we design the Frequency Selective Module (FSM) for moiréd sRGB features, leveraging a learnable band stop filter to mitigate moiré patterns in the frequency domain. Subsequent to the pre-demoiréing operations, we develop the sRGB Guided Image Signal Processor (RGISP) to integrate color information from the coarsely demoiréd sRGB features with the coarsely demoiréd RAW features during the ISP stage, thereby facilitating the color recovery process. Ultimately, for image reconstruction, we utilize multiple Residual Swin Transformer Blocks (RSTBs) [13] and convolutions to accomplish global tone mapping and detail refinement.
@@
    We propose to exploit image demoiréing in both RAW and sRGB domains. The incorporation of RAW data enhances the moiré pattern removal process, while the inclusion of sRGB data reduces the difficulty of RAW-to-sRGB inversion and aids in color restoration of images.
@@
    We develop SCDM with specifically designed modules embeded in the skip-connections to perform efficient and effective demoiréing on the RAW and sRGB features.
@@
    RGISP is designed for RAW-to-sRGB conversion, incorporating color information from coarsely demoiréd sRGB feature with demoiréd RAW features during the ISP stage, facilitating the process of color deviation correction.
@@
    Extensive experiments are conducted on the RAW demoiréing dataset. Our RRID surpasses state-of-the-arts methods in both qualitative and quantitative evaluations.
@@ 2.1 Image and Video Demoiréing
To mitigate moiré patterns resulting from frequency aliasing, numerous learning-based image demoiréing methods [2, 7, 32, 16, 22, 17, 21] and video demoiréing methods [3, 18, 25, 29] have been introduced. DMCNN [20] introduces the first real-world dataset for image demoiréing along with a multi-resolution convolutional neural network. MopNet [6] devises a multi-scale aggregated network guided by edges and attentive to pattern attributes for moiré pattern removal. WDNet [15] employs wavelet transformation to decompose images with moiré patterns into separate frequency bands and establishes a dual-branch network to restore both close-range and far-range information. FHDMi [7] formulates a two-stage approach for simultaneous removal of substantial moiré patterns while preserving image details. ESDNet [26] explores a lightweight model aimed at ultra-high-definition image demoiréing. VDMoiré [3] conducts the first video demoiréing dataset and presents a baseline video demoiréing model that utilizes three consecutive frames. The aforementioned methods all operate in the sRGB domain. However, it has been observed that moiré patterns are less prominent in the RAW domain, which aids in the process of moiré pattern removal. Therefore, some researchers advocate performing image or video demoiréing in the RAW domain. RDNet [27] introduces the first image demoiréing dataset with Raw data and conducts the demoiréing process in the RAW domain. RawVDemoiré [29] proposes a temporal alignment method for Raw video demoiréing. To effectively remove moiré patterns in the RAW domain, the model must possess the simultaneous capabilities of demoiréing and domain conversion. However, previous methods encounter challenges in color recovery due to the limited availability of relevant information in the ISP. Consequently, we propose an image demoiréing network that utilizes both sRGB and Raw data, explicitly addressing the aforementioned issues.
Figure 2: The overview and detailed structures of our proposed RRID for image demoiréing.
@@ 2.2 Learning-Based RAW Image Processing
Raw pixels offer the potential for leveraging additional information due to their broader bit width, inherently containing a greater wealth of data. Early research in Raw image processing primarily centers on integrating demosaicking and denoising techniques [10, 5, 11, 14]. RTF [10] demonstrates a machine learning approach to the demosaicing problem, providing versatility by being applicable to various tiled CFA patterns. SGNet [14] introduces a self-guidance network that incorporates density-map guidance and green-channel guidance for the combined tasks of demosaicing and denoising. Recently, numerous low-level vision tasks endorse the utilization of Raw data, encompassing low-light enhancement [8, 4, 9], super-resolution [24, 28], and reflection removal [12, 19]. The majority of the previously mentioned Raw image processing methods only take Raw images as input and generate restored sRGB images as output, indicating that the entire model is responsible for both the restoration and ISP tasks. CR3Net [19] presents a cascaded network that harnesses both the RGB data and their corresponding Raw versions for image reflection removal. Within CR3Net, a module is formulated to emulate point-wise mappings within the ISP, translating Raw features into the sRGB domain. To further mitigate the color casting problem, our proposed method incorporates guidance from sRGB information with the Raw information within the ISP process.
@@ 3 Methodology
The purpose of image demoiréing is to restore the original clean image by removing moiré patterns and correcting color deviations from an contaminated image. Originally caused by frequency interference between the camera’s CFA and the screen’s LCD subpixel, moiré patterns are further deteriorate by the non-linear processes in ISP. Therefore, introducing raw data for image demoiréing may offer greater efficiency. However, relying solely on raw data for image demoiréing is insufficient in mitigating color cast due to the absence of essential information required for color correction by the Image Signal Processor (ISP). Subsequently, we propose to concurrently utilize both RAW and sRGB data, which is readily accessible in both smartphones and digital cameras, for image demoiréing.
Fig. 2 presents an overview of our proposed RRID for image demoiréing. The input of the RGB stage is an sRGB image $\mathbf{I}_{rgb}\in\mathbb{R}^{H\times W\times 3}$, where $H\times W$ represents the spatial resolution. Given an input RAW image, it is first packed into the 4-channel RGGB format $\mathbf{I}_{raw}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 4}$, which serves as the input for the RAW stage. RRID is designed to generate both the final output sRGB image ($\mathbf{\hat{Y}}_{rgb}$) and a RAW image ($\mathbf{\hat{Y}}_{raw}$) for supervision.
Concretely, shallow features $\mathbf{F}_{raw}$ and $\mathbf{F}_{rgb}$ are initially extracted using a convolutional layer and the Dilated Channel Attention Block (DCAB). It is worth noting that a down-sampling layer is applied to the sRGB data to ensure the uniformity of feature shapes. Then, we apply SCDM to acquire pre-demoiréd features $\mathbf{D}_{raw}$ and $\mathbf{D}_{rgb}$ in RAW stage and RGB stage, respectively. GFM is introduced in the skip-connection of SCDM for raw features, enabling adaptive differentiation between texture details and moiré patterns through feature gating. Similarly, FSM is utilized to to mitigate moiré patterns in the frequency domain. More details will be given in Sec. 3.1. Subsequently, RGISP is designed to realize a raw-to-rgb conversion, acquiring $\mathbf{D}_{out}$. RGISP incorporates color information from coarsely demoiréd sRGB features $\mathbf{D}_{rgb}$ in a cross-attention mechanism, which will be introduced in Sec. 3.2. For the final image reconstruction, 4 RSTBs [13] are utilized for their advantages in building long-range dependencies for global tone mapping and detail refinement.
@@ 3.1 Design of SCDM
We propose SCDM, a multi-scale architecture designed to remove moiré patterns from shallow RAW features $F_{raw}$ and shallow sRGB features $F_{rgb}$, as shown in Fig. 2(a). Unlike the traditional vanilla U-Net, the SCDM is constructed upon DCAB, facilitating the encoding and decoding of informative features within the respective layers. The features are downsampled twice in the spatial dimension during the encoding process and then upsampled to the original resolution of the input features during decoding. To emphasize the contextual information, we enhance the learned features from the corresponding layers of the encoder with additional modules at skip connections, that is GFM for RAW and FSM for sRGB. The introduction of carefully designed demoiréing modules in skip-connections is also intended to enhance model efficiency. Then, the pre-demoiréd features $\mathbf{D}_{raw}$ and $\mathbf{D}_{rgb}$ are acquired.
DCAB. As the essential building block of encoder-decoders for realizing RAW demoiréing and the RGB demoiréing, the detailed structure of DCABs is shown in Fig. 2(b). DCAB is designed using a series of dilated convolution layers followed by ReLU activation and the channel attention mechanism. The receptive field at each block is expanded with the dilated convolution kernels, facilitating the demoiréing operations. Furthermore, by incorporating the channel attention mechanism [31], DCAB adaptively rescales features according to interdependencies among channels. Also, shortcuts are introduced to establish connections between the input and output of DCAB, enabling the block to learn the residual features.
GFM. Embeded in the skip-connection of SCDM for raw features, GFM is introduced to enable adaptive differentiation between texture details and moiré patterns through feature gating. For efficiency, we employ point-wise and depth-wise convolutions to aggregate channel and local content information, respectively. Then, the features are chunked along the channel dimension to $\mathbf{F}_{gate}$ and $\mathbf{F}_{content}$. A point-wise multiplication is applied on $\mathbf{F}_{content}$ and $\mathbf{F}_{gate}$ after a $\mathrm{GELU}$ activation. During feature gating, we expect that the original image content is adaptively selected and merged along both spatial and channel dimensions. The detailed structure of the GFM is shown in Fig. 2(c). The operations can be formulated as:
$\mathbf{F}_{gate},\mathbf{F}_{content}=\mathrm{DConv}(\mathrm{PConv}(\mathbf{F}_{GFM\_in})),$ (1)
@@
where $\mathrm{DConv}$ and $\mathrm{PConv}$ represent depth-wise convolution and point-wise convolution, respectively. $\odot$ denotes the element-wise multiplication.

### 15 Mar 2024 v2

# Image Demoiréing in RAW and sRGB Domains

@@ Abstract
Moiré patterns frequently appear when capturing screens with smartphones or cameras, potentially compromising image quality. Previous studies suggest that moiré pattern elimination in the RAW domain offers greater effectiveness compared to demoiréing in the sRGB domain. Nevertheless, relying solely on RAW data for image demoiréing is insufficient in mitigating the color cast due to the absence of essential information required for the color correction by the image signal processor (ISP). In this paper, we propose to jointly utilize both RAW and sRGB data for image demoiréing (RRID), which are readily accessible in modern smartphones and DSLR cameras. We develop Skip-Connection-based Demoiréing Module (SCDM) with Gated Feedback Module (GFM) and Frequency Selection Module (FSM) embedded in skip-connections for the efficient and effective demoiréing of RAW and sRGB features, respectively. Subsequently, we design a RGB Guided ISP (RGISP) to learn a device-dependent ISP, assisting the process of color recovery. Extensive experiments demonstrate that our RRID outperforms state-of-the-art approaches, in terms of the performance in moiré pattern removal and color cast correction by 0.62dB in PSNR and 0.003 in SSIM.
Keywords:
Image demoiréing RAW domain Image restoration
@@ 1 Introduction
With the increasing prevalence of smartphone cameras, it is common to record on-screen content using smartphone cameras for convenience. However, this type of photography often results in the presence of moiré patterns in the captured images. The appearance of moiré patterns is caused by the spatial frequency aliasing between the camera’s color filter array (CFA) and the screen’s LCD subpixel [21], resulting in an unsatisfactory visual experience.
@@
Moiré patterns present challenges to be eliminated due to their different scales, indistinct shapes, diverse colors, and varying frequencies. Nowadays, many learning-based image demoiréing approaches [20, 1, 5, 6, 14, 31, 15, 22, 26, 17, 21] have been proposed. Among the various methods employed, a commonly used approach is utilizing a multi-scale architecture to remove various sizes of moiré patterns [20, 1, 5, 26, 17]. Besides, some works advocate the leverage of frequency components and delicately design frequency-based demoiréing networks [6, 14, 31, 21].
@@
The aforementioned methods are developed to demoiré in the sRGB domain, as depicted in Fig. 1(a). While this scheme seems to be the most direct approach, its effectiveness is limited in removing moiré patterns from complex scenes. The reason for this limitation probably lies in the fact that the nonlinear operations in ISP such as demosaicing deteriorate the moiré patterns originally in the RAW domain. Many schemes consequently suggest that eliminating moiré patterns in the RAW domain is more effective than performing demoiréing in the sRGB domain [28, 27]. Due to the easy accessibility of RAW data in modern smartphones and DSLR cameras, image demoiréing in RAW domain can be both feasible and advantageous. RDNet [28] is the pioneering study that investigates image demoiréing in the RAW domain. They explore the demoiréing of RAW images using an encoder-decoder and class-specific learning approach, followed by the utilization of a pre-trained ISP for the RAW-to-sRGB conversion, as depicted in Fig. 1(b). Nevertheless, relying exclusively on RAW data can lead to color cast due to the uncertainty during RAW-to-sRGB conversion.
@@
In order to correct the color cast and make the best use of RAW and sRGB information for the moiré pattern removal, we propose to jointly utilize both RAW and sRGB data for image demoiréing (RRID), as demonstrated in Fig. 1(c). We advocate that introducing paired RAW and sRGB data is advantageous to the moiré pattern removal due to the following reasons: (1) RAW pixels provide more information compared to sRGB pixels since they are typically 12 or 14 bits. (2) Moiré patterns in the RAW domain are less apparent because they are not further affected by nonlinear operations in the ISP. (3) The paired RAW and sRGB data allows the model to learn a device-dependent ISP to aid the process of color recovery. In RRID, we adopt a multi-scale architecture in the Skip-Connection-based Demoiréing Module (SCDM), to effectively eliminate moiré patterns of different sizes. In order to effectively perform demoiréing on RAW and sRGB features, the Gated Feedback Module (GFM) and Frequency Selective Module (FSM) are integrated into skip-connections within SCDM, as opposed to brutally injecting them into the multi-scale layers. Specifically, for RAW features demoiréing, we introduce the GFM into the skip connection, enabling adaptive differentiation between texture details and moiré patterns through feature gating. Moreover, we design the FSM for the moiré sRGB features, leveraging a learnable band stop filter to mitigate moiré patterns in the frequency domain. Subsequent to the pre-demoiréing operations, we develop the RGB Guided Image Signal Processor (RGISP) to integrate color information from the demoiréd sRGB features with the demoiréd RAW features and learn a device-dependent ISP, facilitating the color recovery process. Ultimately, for the image reconstruction, we utilize multiple Residual Swin Transformer Blocks (RSTBs) [12] and convolutions to accomplish global tone mapping and detail refinement.
@@
    We propose to exploit image demoiréing in both RAW and sRGB domains. The incorporation of RAW data enhances the moiré pattern removal process, while the inclusion of sRGB data facilitates the RAW-to-sRGB conversion and aids in color restoration of images.
@@
    We develop SCDM with specifically designed modules, GFM and FSM, embeded in the skip-connections to perform efficient and effective demoiréing on the RAW and sRGB features.
@@
    RGISP is designed for the RAW-to-sRGB conversion, incorporating color information from the demoiréd sRGB features with demoiréd RAW features and learn a device-dependent ISP, aiding the color deviation correction.
@@
    Our RRID surpasses state-of-the-arts methods in both qualitative and quantitative evaluations.
@@ 2.1 Image and Video Demoiréing
The purpose of image demoiréing is to restore the original clean image by removing moiré patterns and correcting color deviations from an contaminated image. Many learning-based image demoiréing [1, 6, 31, 15, 22, 17, 21] and video demoiréing methods [2, 18, 25, 27] have been introduced to mitigate the moiré patterns. DMCNN [20] constructs the first real-world dataset for image demoiréing and proposes a multi-resolution neural network to handle moiré patterns across various scales. MopNet [5] devises a multi-scale aggregated network that leverages the edge guidance and pattern attributes for moiré pattern removal. WDNet [14] employs wavelet transformation to decompose images with moiré patterns into separate frequency bands and establishes a dual-branch network to restore both close-range and far-range information. FHDMi [6] formulates a two-stage approach for simultaneous removal of substantial moiré patterns while preserving image details. ESDNet [26] explores a lightweight model aimed at ultra-high-definition image demoiréing. VDMoiré [2] collects the first video demoiréing dataset and presents a baseline video demoiréing model with a relation-based temporal consistency loss. The aforementioned methods all operate in the sRGB domain. Since it has been observed that moiré patterns are less prominent in the RAW domain, some schemes advocate performing image or video demoiréing in the RAW domain. RDNet [28] introduces the first image demoiréing dataset with RAW data and conducts the demoiréing process in the RAW domain. RawVDemoiré [27] proposes a temporal alignment method for RAW video demoiréing. To remove moiré patterns in the RAW domain, the model have to simultaneously possess the capabilities of demoiréing and and RAW-to-sRGB conversion. However, previous methods encounter challenges in color recovery due to the uncertainty of RAW-to-sRGB conversion. Consequently, we propose an image demoiréing network that utilizes paired RAW and sRGB data, facilitating the color recovery process in image demoiréing.
@@ 2.2 Learning-Based RAW Image Processing
RAW pixels offer the potential for leveraging additional information due to their broader bit depth, inherently containing a great wealth of data. Early research in RAW image processing primarily centers on integrating demosaicing and denoising techniques [9, 4, 10, 13]. RTF [9] devices a machine learning approach to the demosaicing and denoising on RAW inputs. SGNet [13] introduces a self-guidance network that leverages the inherent density-map guidance and green-channel guidance for demosaicing and denoising on the RAW images. Recently, some low-level vision tasks endorse the utilization of RAW data, including low-light enhancement [7, 3, 8], super-resolution [24, 29], and reflection removal [11, 19]. The majority of the aforementioned RAW image processing methods only take RAW images as input and generate restored sRGB images as output, indicating that the entire model is responsible for both the restoration and ISP tasks. CR3Net [19] presents a cascaded network that harnesses both the RGB data and their corresponding RAW versions for image reflection removal. Within CR3Net, a module is formulated to emulate point-wise mappings within the ISP, converting the features from RAW to sRGB domain. To further mitigate the color casting problem, our proposed method jointly utilize RAW and sRGB data to learn a device-dependent ISP for image demoiréing.
@@ 3 Methodology
Figure 2: The overview and detailed structures of our proposed RRID.
Originally caused by frequency interference between the camera’s CFA and the screen’s LCD subpixel, moiré patterns are further deteriorated by the nonlinear processes in ISP. Therefore, introducing RAW data for image demoiréing may offer greater effectiveness. However, relying solely on RAW data for image demoiréing is insufficient in mitigating color cast due to the uncertainty during the RAW-to-sRGB conversion. Consequently, in this work, we propose to utilize paired RAW and sRGB data, which are readily accessible in modern smartphones and DSLR cameras (e.g., iPhone 15 Pro, HUAWEI P60 Pro). The RAW-sRGB pairs essentially allows the model to learn a device-dependent ISP, aiding the color recovery process. The potential applications in real-world scenarios can be described as follows. When demoiréing function is enabled, the in-device ISP generated sRGB image will not be directly outputted. Instead, together with its corresponding RAW version, will be processed by our RRID to generate a new sRGB image with much suppressed moiré patterns. Afterwards, the RAW image could be discarded, depending on settings of the camera.
More specifically, Fig. 2 presents an overview of our proposed RRID for image demoiréing with the paired RAW and sRGB data. As for the RGB branch, the input is an sRGB image $\mathbf{I}_{rgb}\in\mathbb{R}^{H\times W\times 3}$, where $H\times W$ represents the spatial resolution. Given an input RAW image, it is first packed into the 4-channel RGGB format $\mathbf{I}_{raw}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 4}$, which serves as the input for the RAW branch. RRID is designed to generate the final output sRGB image ($\mathbf{\hat{Y}}_{rgb}$).
Concretely, shallow features $\mathbf{F}_{raw}$ and $\mathbf{F}_{rgb}$ are respectively extracted from the input data $\mathbf{I}_{raw}$ and $\mathbf{I}_{rgb}$ using a convolutional layer and the Dilated Channel Attention Block (DCAB). It is worth noting that a down-sampling layer is applied to the sRGB data to ensure the uniformity of feature shapes. Then, we apply SCDM to acquire pre-demoiréd features $\mathbf{D}_{raw}$ and $\mathbf{D}_{rgb}$ in RAW branch and RGB branch, respectively. GFM is introduced in the skip-connection of SCDM for RAW features, enabling adaptive differentiation between texture details and moiré patterns through feature gating. Similarly, FSM is utilized to to mitigate moiré patterns in the frequency domain. More details will be given in Sec. 3.1. Subsequently, RGISP is designed to realize a RAW-to-sRGB conversion, generating $\mathbf{D}_{out}$. RGISP incorporates color information from coarsely demoiréd sRGB features $\mathbf{D}_{rgb}$ in a cross-attention mechanism, which will be introduced in Sec. 3.2. For the final image reconstruction, 4 RSTBs [12] are utilized to generate sRGB output $\mathbf{\hat{Y}}_{rgb}$ for their advantages in building long-range dependencies for global tone mapping and detail refinement.
@@ 3.1 Design of SCDM
We propose SCDM, a multi-scale architecture designed to remove moiré patterns from shallow RAW features $F_{raw}$ and shallow sRGB features $F_{rgb}$, as shown in Fig. 2(a). Unlike the vanilla U-Net, SCDM is constructed upon DCAB, facilitating the encoding and decoding of informative features within the respective layers. The features are downsampled twice in the spatial dimension during the encoding process and then upsampled to the original resolution of the input features during decoding. To emphasize the contextual information, we enhance the learned features from the corresponding layers of the encoder with additional modules at skip connections, that is GFM for RAW and FSM for sRGB. The introduction of GFM and FSM in skip-connections is also intended to enhance model efficiency. Then, we can obtain the pre-demoiréd features $\mathbf{D}_{raw}$ and $\mathbf{D}_{rgb}$.
DCAB. DCAB is the essential building block of encoder-decoders for performing RAW demoiréing and the RGB demoiréing, as shown in Fig. 2(b). DCAB is designed using a series of dilated convolution layers followed by ReLU activation and the channel attention mechanism. The receptive field at each block is expanded with the dilated convolution kernels, facilitating the demoiréing operations. Furthermore, by incorporating the channel attention mechanism, DCAB adaptively rescales features according to interdependencies among channels. Also, shortcuts are introduced to establish connections between the input and output of DCAB, enabling the block to learn the residual features.
GFM. Embeded in the skip-connection of SCDM for RAW features, GFM is introduced to enable adaptive differentiation between the texture and moiré patterns through feature gating. The detailed structure of the GFM is shown in Fig. 2(c). For efficiency, we employ point-wise and depth-wise convolutions to aggregate channel and local content information, respectively. Then, the features are chunked along the channel dimension to generate $\mathbf{F}_{gate}$ and $\mathbf{F}_{content}$. A point-wise multiplication is applied on $\mathbf{F}_{content}$ and $\mathbf{F}_{gate}$ after a $\mathrm{GELU}$ activation. During feature gating, we expect that the original image content is adaptively selected and merged along both spatial and channel dimensions. The operations can be formulated as:
$\{\mathbf{F}_{gate},\mathbf{F}_{content}\}=\mathrm{DConv}(\mathrm{PConv}(\mathbf{F}_{GFM\_in})),$ (1)
@@
where $\mathrm{DConv}$ and $\mathrm{PConv}$ represent depth-wise convolution and point-wise convolution, respectively. Here, $\odot$ denotes the element-wise multiplication.
FSM. For sRGB demoiréing, we leverage a learnable band stop filter [31] to mitigate moiré patterns in the frequency domain. Considering the formation of moiré patterns is caused by the spatial frequency aliasing between the camera CFA and screen’s LCD subpixel, demoiréing in the frequency domain becomes a favorable approach. However, the frequency spectrum of the moiré pattern is often mixed with that of the original contents, thereby making it challenging to disentanglement. Therefore, we propose to perform the disentanglement within small patches, where the moiré patterns tend to be more pronounced in a specific range of frequency bands, facilitating the removal of moiré patterns. Inspired by Zheng et al. [31], we utilize a band stop filter to amplify certain frequencies and suppress others using Block DCT. However, obtaining the frequency prior to separate moiré patterns from normal image textures is challenging and time-consuming. To address this issue, a set of convolution layers are employed as the learnable band stop filter $\mathrm{B}(\cdot)$ to attenuate the specific frequencies of moiré patterns while preserving the original image contents. The process of the FSM can be represented as:
where $\mathrm{DCT}$ and $\mathrm{IDCT}$ denote the process of Block DCT and the corresponding inverse process. The block size is set to $8\times 8$ for $\mathrm{DCT}$ and $\mathrm{IDCT}$.
3.2 Design of RGISP
RGISP is proposed to transform pre-demoiréd RAW features $\mathbf{D}_{raw}$ into sRGB domain $\mathbf{D}_{out}$. With the assistance of pre-demoiréd sRGB features $\mathbf{D}_{rgb}$, a device-dependent ISP can be learned in RGISP. Matrix transformation is commonly employed in the traditional ISP pipelines [16]. It enables the enhancement or conversion of image colors to another color space through a channel-wise matrix transformation, which is facilitated by globally shared settings such as environmental illumination and color space specifications [8]. Building upon this principle, we introduce matrix transformation as a means to perform color transformation, as illustrated in Fig. 3. The design of this matrix transformation is motivated by recent advancements in transposed cross-attention.
Figure 3: The structure of RGISP.
Given $\mathbf{D}_{raw}\in\mathbb{R}^{C\times\frac{H}{2}\times\frac{W}{2}}$, the vectors of query $\mathbf{Q}\in\mathbb{R}^{C\times\frac{H}{2}\frac{W}{2}}$, key $\mathbf{K}\in\mathbb{R}^{\frac{H}{2}\frac{W}{2}\times C}$ are generated through the projection with a 1 $\times$ 1 convolutional layer $\mathrm{Conv}$ and a flatten operation $\mathrm{Flatten}$, where $C$ is the intermediate channel number. Similarily, value $\mathbf{V}\in\mathbb{R}^{C\times\frac{H}{2}\frac{W}{2}}$ can be produced by operations on $\mathbf{D}_{rgb}\in\mathbb{R}^{C\times\frac{H}{2}\times\frac{W}{2}}$ as:
$\{\mathbf{Q},\mathbf{K}\}=\mathrm{Flattern}(\mathrm{Conv}(\mathbf{D}_{raw})),$ (4)
$\mathbf{V}=\mathrm{Flatten}(\mathrm{Conv}(\mathbf{D}_{rgb})).$ (5)
where a scaling coefficient $\lambda$ is applied for numerical stability. Subsequently, the vector $\mathbf{V}$ is transformed by the matrix $\mathbf{M}$, performing color space conversion in feature-level. The output feature after color transformation can be obtained by $\mathbf{D}_{out}=\mathbf{M}\cdot\mathbf{V}$. As a complement to the global matrix transformation, we use a depth-wise convolution and a point-wise convolution to refine the local details, generating the output in sRGB domain $\mathbf{D}_{out}$.
3.3 Loss Function
To effectively remove the moiré patterns and adjust the color deviations, we introduce supervision in both RAW and sRGB domains. The overall training objective can be expressed as:
where ${\mathbf{Y}}_{raw}$ and ${\mathbf{Y}}_{rgb}$ are the ground-truth RAW and sRGB images, respectively. We empirically set $\alpha$=0.5.
4 Experiments
4.1 Experimental Setup
Dataset. The experiments are conducted on the RAW image demoiréing dataset published in TMM22 [28], which contains recaptured scenes of natural images, documents, and webpages. In TMM22, the dataset consists of 540 RAW and sRGB image pairs with ground-truth for training, and 408 pairs for testing. To facilitate the training and comparison process, patches of size 256$\times$256 and 512$\times$512 are cropped respectively for the training and testing sets. The moiré images are captured using four different smartphone cameras and three display screens of varying sizes. In addition, we utilize an image demoiréing dataset FHIMi [6] to verify the generalization ability of our method. FHDMi contains 9981 sRGB image pairs for training and 2019 for testing, featuring diverse and intricate moiré patterns. Notably, images in FHDMi have a higher resolution of 1920$\times$1080, compared to TMM22.
Training Details. We train RRID using the AdamW optimizer with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999. A multistep learning rate schedule is employed, with the learning rate initialized at 2×$\times 10^{-4}$. The RRID model is trained for 500 epochs using a batch size of 80 on 4 NVIDIA RTX 3090 GPUs.
4.2 Quantitative Results
We compare our approach with several image demoiréing methods in sRGB domain, including DMCNN [20], WDNet [14], ESDNet [26], and a RAW image demoiréing method RDNet [28]. We further add comparison with a low-light enhancement method DNF [8] with RAW input, a video demoiréing method RawVDmoiré [27] with RAW input, and a de-reflection method CR3Net [19] with paired RAW-sRGB data. To evaluate the performance of our proposed RRID, we adopt the following three standard metrics to assess pixel-wise accuracy and the perceptual quality: PSNR, SSIM [23], and LPIPS [30]. Additionally, the number of parameters, FLOPs, and inference time are introduced to measure the model complexity. To ensure a fair comparison, we modify the inputs of the sRGB-based models to align with the dataset requirements. Subsequently, all models are fine-tuned using the default settings as provided in their papers on TMM22 dataset. We select the better results from the pretrained and retrained models for comparison, which provides an advantage over the competing methods.